1. Gorgeousleny with an income of P150 would like to buy two goods. Let us say juice and bread. Price of juice is P12 and price of bread is P6. If gorgeousleny will buy the ist 5 juices she is entitled to a 50% off. Draw the graph.

After a week gorgeousleny went back to the store and will still buy her two favorite goods with the same income of P150 and price of bread and juice is constant. but this time the promo of 50% off is applied after the ist 5 juices. Draw the graph. .

a. ) Measure the qty. of bread on the vertical axis and the qty. of juice on the horizontal axis. Label both graph with its corresponding points, pt. A for bread and point B for juice. the combination of bread and juice label it as point C.

b.) For both graph, What is the slope from point C to point B? ; What is the slope from from point A to point C?  



1
Expert's answer
2021-09-20T11:04:30-0400

Budget Constraint:

Price of Bread"\times" Quantity of Bread + Price of Juice"\times" Quantity of Juice = Income

"6\times15 + 12\times5= 90+ 60 = 150"

New price of juice 6

New Budget Constraint:

"6\times Bread + 12\times Juice = 150"

If she purchase bread and no juice:

"6\times Bread + 12\times 0 = 150"

Bread= 25


Slope of Budget Constraint is price ratio. Therefore,

(a) 


Black is old combination bundle-

Green is new combination bundle- 

(b)

Slope of old budget line:"\frac{12}{6} = 2"

Slope of new budget line:"\frac{6}{6} = 1"